# FEM-based Neural Networks for Solving Incompressible Fluid Flows and Related Inverse Problems

### v1

@@ 1 Introduction
Among the numerous approaches to utilize the power of neural networks for solving differential equations, in particular so-called physics-informed neural networks (PINNs) [6] gained a lot of attention in the last years. PINNs embed the residual of the respective differential equations into the loss function of a neural network and compute the required derivatives exactly by automatic differentiation [7]. Hence, this approach is often considered to be able to avoid direct discretization; in particular, it is clearly a mesh-free approach. Since training exploits the underlying physics, encoded in the loss function, the PINN approach is able to solve a differential equation without the need for simulated or measured data. However, as higher order derivatives of the network in the loss function lead to a high learning complexity, and additional loss terms that incorporate boundary conditions lead to a multi-objective optimization problem (at least in the vanilla PINN approach), training PINNs is known to be potentially challenging; see, e.g., [8]. For further literature dealing with PINNs, extensions thereof, or similar techniques such as the Deep Ritz method, the Deep Nitzsche method, and weak adversarial networks — all of these methods have in common that they do not explicitly apply discretization — we refer to the aforementioned survey papers.
@@
In the following, we focus our literature overview on physics-informed neural networks that make explicit use of discretization, at least to a certain degree. The so-called variational PINNs (vPINNs), a Petrov–Galerkin-like approach, in which a neural network serves as trial function that is tested with a set of finitely many test functions, are introduced in [9]. In the VarNet approach [10], specifically finite element functions are used as test functions, and in $hp$-vPINNs [11] essentially $hp$-finite element functions are used as test functions. In contrast, the Galerkin neural network approach [12] extends the classical Galerkin procedure, that is usually bound to trial and test functions coming from linear spaces, to trial and test functions both coming from non-linear sets of functions parameterized by neural networks. PhyGeoNet [13] maps the underlying problem to a rectangular reference domain and then employs PINNs loss evaluated on a rectangular grid of points together with convolutional neural networks. Extending this idea to less structured meshes, physics-informed graph neural Galerkin networks [14] combine graph convolutional NNs with a Galerkin finite element method (FEM) discretization. In [15], finite element discretization is considered and the proposed network architecture mimics subsequent refinement steps of the underlying mesh; moreover, a block Jacobi preconditioner is applied. Recently, the use of PINNs has also been proposed to speed up certain element-local computations within a finite element method [16], and neural networks have been used to correct coarse level FEM-solutions on a finer level [17, 18]. At the end of this overview, let us briefly comment also on two methods with a similar title that are only slightly related to the present context. In [19], the differential equation is discretized using classical FEM, while a neural network is used to parameterize an unknown parameter function in the differential equation. In [20], the finite element method is cast into a neural network.
@@
Finally, [21, 22] combine neural networks and finite elements directly: the parametric differential equation under consideration is discretized a priori by finite elements, and a neural network is trained to map the parameters (of the differential equation) to the coefficients of the corresponding discrete solution (w.r.t. the chosen finite element basis). The loss function is given by the residual of the FEM-discretized differential equation, making this approach data-free and physics-informed. In [21] the applicability to 1D-convection-diffusion problems, as well as to a truss problem, and in [22] to further elasticity problems, such as the Euler–Bernoulli beam or the Reissner–Mindlin plate, has been demonstrated. For a discussion of the advantages of this approach compared to PINNs, in particular the simplified handling of boundary conditions or irregularly shaped domains, we refer to [21]. While [21, 22] are concerned with numerical parameters, an extension to a still fully discretized, but more operator-learning-like scenario has been presented in [23] in the context of a time-dependent linear heat equation.
@@
In our opinion, combining the mathematically proven and quantified ability of FEMs to solve differential equations with the strong approximation capabilities of neural networks to deal with parameter dependencies, should be considered as a simple, but highly appealing idea. In the present work, we thus extend this approach in two main directions:
@@
    First, we consider stationary 2D Stokes and Navier–Stokes equations, i.e., equations of type and characteristics that, given their inherent saddle-point structure, are even for classical FEM more challenging than those previously treated with this method.
@@
    Second, we modify the physics-informed loss function by adding a preconditioner. In conjunction with L-BFGS, this significantly speeds up training and finally improves accuracy of the obtained solutions.
@@
In Sect. 2 we introduce the proposed method; first, we describe the underlying setting and the idea in general, and then we discuss this approach in more detail on behalf of our two concrete model problems in Sects. 2.3 and 2.4. Numerical examples illustrating the advantages of our method are given in Sect. 3, and a prototypical application to probabilistic inverse problems is presented in Sect. 4. We conclude the paper with a brief summary and outlook on future work in Sect. 5. Finally, the appendix Sect. A contains the proof of a fact that can be understood as a heuristic motivation for the chosen approach.
@@ 2.1 Problem statement
Let now $V_{h}\subset V$ be a finite dimensional subspace, e.g., resulting from a conforming finite element discretization of $V$; the index $h$ indicates the discretization parameter, e.g., the mesh size used for finite element discretization. The so-called Galerkin discretization of our problem (2) is given by
@@ 2.2 Proposed Method
Since (4) is a possibly nonlinear system of equations, usually consisting of a high number of variables, solving (4) for several choices of $\lambda$ may become costly. To deal with this issue, we propose a method that makes use of the strong approximation capabilities of neural networks. More precisely, we want to approximate the map $\lambda\mapsto{\bf u}_{h}(\lambda)$ solving (4) by a neural network ${\bf u}_{h,\textnormal{NN}}^{\theta}\!\colon\mathbb{R}^{p}\rightarrow\mathbb{R}^{N_{h}}$; hereby, $\theta$ denotes the network parameters of the neural network. Training this neural network — or in other words: finding the right parameters $\theta$ that make ${\bf u}_{h,\textnormal{NN}}^{\theta}$ a hopefully good approximation of $\lambda\mapsto{\bf u}_{h}(\lambda)$ — requires to determine a suitable so-called loss function. Previous work [21] has suggested to use the $\ell^{2}$-norm of the discrete equation residual, i.e. to determine $\theta=\operatorname{argmin}_{\vartheta}\mathcal{L}$ of
@@
with ${\bf P}^{L}_{\lambda,h},{\bf P}^{R}_{\lambda,h}\in\mathbb{R}^{N_{h}\times N_{h}}$ being the left, and respectively right, preconditioners. The desired ${\bf u}_{h,\textnormal{NN}}^{\theta}(\lambda)$ can then be reobtained via ${\bf u}_{h,\textnormal{NN}}^{\theta}(\lambda)={\bf P}^{R}_{\lambda,h}{\bf\tilde{u}}_{h,\textnormal{NN}}^{\theta}(\lambda)$. Of course, such a preconditioner must be chosen carefully taking into account the problem characteristics.
@@
To optimize the loss function w.r.t. the neural network parameters $\theta$, usually variants of the gradient descent method, such as the Adam optimizer [27], or quasi-Newton methods, such as the limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) algorithm [28] are employed. These methods require the gradient of the loss function with respect to the network parameters, which is calculated by using automatic differentiation (AD) [7]; thus all computations in the FEM Module, such as, e.g., setting up stiffness and/or mass matrices etc., must be implemented automatically differentiable. In this paper we will solely employ L-BFGS. Like most optimizers popular in deep learning, L-BFGS only requires gradients of the objective function, i.e., first-order information, and thus avoids the expensive computation of second-order derivatives by AD, which would be necessary for second-order methods as, e.g., Newton’s method. Nevertheless, as a quasi-Newton method, L-BFGS lies somewhere between first- and second-order methods as it is able to exploit curvature information by successively building an approximation to the Hessian matrix during the iteration. The associated hopes for faster convergence have been confirmed, e.g., in the case of classic PINNs [6].
@@
Hence, we hope to improve the conditioning of the Hessian of $\mathcal{L}_{{\bf P}_{h}}$ at some minimizer $\theta$ by choosing ${\bf P}^{L}_{\lambda,h}$, ${\bf P}^{L}_{\lambda,h}$ in such a way that the condition number of $\mathbf{P}^{L}_{\lambda,h}\cdot[D_{\bf u}{\bf R}_{\lambda,h}](\mathbf{P}^{R}_{\lambda,h}{\bf\tilde{u}}_{h,\textnormal{NN}}^{\theta}(\lambda))\cdot\mathbf{P}^{R}_{\lambda,h}$ becomes relatively small. Even if these considerations by no means constitute a strict proof, we consider them to be a conclusive motivation for our approach, and our numerical observations in Sect. 3 seem to support this.
@@ 2.3 Example 1: Stokes equations in 2D
In the following, we show some more concrete details regarding the realization and implementation of the proposed FEM-based neural network for solving the Stokes flow around an airfoil with a parameterizable angle of attack. We consider the problem in a 2-dimensional domain $\Omega$, where the Stokes problem consists of finding two functions, the velocity $u(\lambda)\!:\Omega\rightarrow\mathbb{R}^{2}$ and the pressure $p(\lambda)\!:\Omega\rightarrow\mathbb{R}$ such that
@@
with functions $f_{\lambda}\!:\Omega\rightarrow\mathbb{R}^{2}$, $g_{\lambda}\!:\Omega\rightarrow\mathbb{R}$ and suitable boundary conditions, which will be introduced precisely below. Equation (7) represents the conservation of momentum, equation (8) the conservation of mass.
The discretized counterpart of system (7)–(8) with Galerkin ansatz space $V_{h}:=X_{h}\times M_{h}$ reads as follows:
@@
with the bilinear forms $a(u_{h},v_{h}):=\int_{\Omega}\nabla u_{h}:\nabla v_{h}$ and $b(v_{h},p_{h}):=-\int_{\Omega}p_{h}\nabla\cdot v_{h}$. The two functionals $f$ and $g$ are defined as $f(v_{h})\coloneqq\int_{\Omega}v_{h}f_{\lambda}$ and $g(q_{h})\coloneqq-\int_{\Omega}q_{h}g_{\lambda}$, respectively. $R_{\lambda}$ is accordingly defined as the residual of equation (9). To obtain a well-posed problem, the discrete problem must satisfy the Ladyzhenskaya–Babus̆ka–Brezzi (LBB) condition [30, 31, 32] and hence, the choice of $X_{h},M_{h}$ must be handled with care. We use Taylor–Hood mixed finite elements, i.e., continuous, piecewise linear ansatz functions for the pressure and continuous, piecewise quadratic ansatz functions for the velocity. With this choice of $V_{h}$, the LBB condition for (9) is satisfied and we obtain a well-posed discrete problem  [33, Lemma 4.23].
@@
Lemma 2.1 in [35] shows that the conditioning of the plain saddle point problem (10) is bounded from below by the conditioning of the discretization of the Laplacian. As such, classical FEM theory [33, Theorem 9.14] shows that for a Lagrange finite element, as we have chosen for the velocity discretization, the conditioning scales with $h^{-2}$. Therefore, it is usually necessary to apply preconditioning to the simple saddle-point problem. In fact, the condition number of the square of the preconditioned stiffness matrix in (12) is bounded independently of the discretization, as shown in Sect. A. There, it is even shown that the number of distinct eigenvalues of this matrix is three. Considering again that BFGS with perfect line search is equivalent to CG for quadratic optimization problems [29], BFGS applied to such an quadratic optimization problem with the squared preconditioned matrix as its Hessian, would converge in at most three iterations [^2] The situation is analogous to the Schur-preconditioned stable FEM solutions of the Stokes problem, when solved with Krylov subspace methods, such as MINRES, cf. [36, p. 292].. Of course, these properties do not directly transfer to our setting, as the problem under consideration is highly non-linear and non-convex due to the presence of a neural network. Nevertheless, it does not seem unreasonable to assume that applying the preconditioner to the linear part of the problem will lead to at least a gradual improvement of the overall properties in this case as well. In fact, our numerical examples in Sect. 3 support this.
@@
Let us now turn to the explicit numerical example, where we solve the Stokes equations (7)-(8) with source terms $f_{\lambda}=0,g_{\lambda}=0$, and non-homogeneous Dirichlet boundary conditions. We define the domain as $\Omega=[0,5]^{2}$ with a NACA 0012 airfoil [37] of length one in the center. A Dirichlet boundary condition dependent on the angle of attack $\lambda$ is used at the inflow boundary $\Omega^{+}:=\{(x,y):x=0\text{ or }y=0\}$, a Neumann boundary condition at the outflow boundary $\Omega^{-}:=\{(x,y):x=5\text{ or }y=5\}$ and a no-slip boundary condition at the airfoil boundary $\Omega^{A}$ (see Figure 2):

### v2

# Preconditioned FEM-based Neural Networks for Solving Incompressible Fluid Flows and Related Inverse Problems

@@ 1 Introduction
Among the numerous approaches to utilize the power of neural networks for solving differential equations, in particular so-called physics-informed neural networks (PINNs), first introduced in [6, 7] and brought to broad attention by [8], gained a lot of attention in the last years. PINNs embed the residual of the respective differential equations into the loss function of a neural network and compute the required derivatives exactly by automatic differentiation [9]. Hence, this approach is often considered to be able to avoid direct discretization; in particular, it is clearly a mesh-free approach. Since training exploits the underlying physics, encoded in the loss function, the PINN approach is able to solve a differential equation without the need for simulated or measured data. However, as higher order derivatives of the network in the loss function lead to a high learning complexity, and additional loss terms that incorporate boundary conditions lead to a multi-objective optimization problem (at least in the vanilla PINN approach), training PINNs is known to be potentially challenging; see, e.g., [10]. For further literature dealing with PINNs, extensions thereof, or similar techniques such as the Deep Ritz method, the Deep Nitsche method, and weak adversarial networks — all of these methods have in common that they do not explicitly apply discretization — we refer to the aforementioned survey papers.
@@
In the following, we focus our literature overview on physics-informed neural networks that make explicit use of discretization, at least to a certain degree. The so-called variational PINNs (vPINNs), a Petrov–Galerkin-like approach, in which a neural network serves as trial function that is tested with a set of finitely many test functions, are introduced in [11]. In the VarNet approach [12], specifically finite element functions are used as test functions, and in $hp$-vPINNs [13] essentially $hp$-finite element functions are used as test functions. In contrast, the Galerkin neural network approach [14] extends the classical Galerkin procedure, that is usually bound to trial and test functions coming from linear spaces, to trial and test functions both coming from non-linear sets of functions parameterized by neural networks. PhyGeoNet [15] maps the underlying problem to a rectangular reference domain and then employs PINNs loss evaluated on a rectangular grid of points together with convolutional neural networks. Extending this idea to less structured meshes, physics-informed graph neural Galerkin networks [16] combine graph convolutional NNs with a Galerkin finite element method (FEM) discretization. In [17], finite element discretization is considered and the proposed network architecture mimics subsequent refinement steps of the underlying mesh; moreover, a block Jacobi preconditioner is applied. In [18] a somehow hybrid combination of PINNs and FEM-discretization is combined with preconditioning and applied to several model problems, including stationary Navier–Stokes equations (with homogeneous boundary conditions): while the network still approximates the non-discretized solution, it is only evaluated at grid points of a Lagrange FEM-discretization, and the loss functions for the training is constituted in terms of the left-preconditioned FEM-residual. Nonlinear problems are treated by iteratively solving linear problems. A model problem very similar to the one under consideration in the present paper is dealt with in [19] using classical PINNs again. A so-called time-stepping-oriented training procedure, inspired by an control-theoretic ansatz to improve condition number of Jacobi matrices, is employed and allows to deal with 3d problems up to Reynolds number 5,000. We see these results as an encouraging sign that investigating the use of classical numerical techniques and “tricks” is promising in the context of physics-informed ML.
Recently, the use of PINNs has also been proposed to speed up certain element-local computations within a finite element method [20], and neural networks have been used to correct coarse level FEM-solutions on a finer level; see, e.g., [21, 22, 23]. Finally, a combination of Isogeometric Analysis with classical PINNs has been proposed in [24]. At the end of this overview, let us briefly comment also on two methods with a similar title that are only slightly related to the present context. In [25], the differential equation is discretized using classical FEM, while a neural network is used to parameterize an unknown parameter function in the differential equation. In [26], the finite element method is cast into a neural network. The domain-decomposition-based (nonlinear) preconditioning of the training of physics-informed neural networks developed in [27] refers to layer-parallel decomposition of the respective neural networks.
@@
Finally, [28, 29] combine neural networks and finite elements directly: the parametric differential equation under consideration is discretized a priori by finite elements, and a neural network is trained to map the parameters (of the differential equation) to the coefficients of the corresponding discrete solution (w.r.t. the chosen finite element basis). The loss function is given by the residual of the FEM-discretized differential equation, making this approach data-free and physics-informed. In [28] the applicability to 1D-convection-diffusion problems, as well as to a truss problem, and in [29] to further elasticity problems, such as the Euler–Bernoulli beam or the Reissner–Mindlin plate, has been demonstrated. For a discussion of the advantages of this approach compared to PINNs, in particular the simplified handling of boundary conditions or irregularly shaped domains, we refer to [28]. While [28, 29] are concerned with numerical parameters, an extension to a still fully discretized, but more operator-learning-like scenario has been presented in [30] in the context of a time-dependent linear heat equation. Among the vast amount of literature on operator learning [31], another popular method to address parametric problems, we mention exemplarily the recent contribution on hypersonic flows [32].
@@
In our opinion, combining the mathematically proven and quantified ability of FEMs to solve differential equations with the strong approximation capabilities of neural networks to deal with parameter dependencies, should be considered as a simple, but highly appealing idea. In the present work, we thus stick to this approach and extend it in two main directions:
@@
    First, we consider stationary 2D Stokes and Navier–Stokes equations, i.e., equations of type and characteristics that, given their inherent saddle-point structure, are even for classical FEM quite challenging.
@@
    Second, we modify the physics-informed loss function by adding a preconditioner, from left and right, directly to the (nonlinear) residual. In conjunction with L-BFGS, this significantly speeds up training and finally improves accuracy of the obtained solutions.
It should be noted that preconditiong from both sides and solving the nonlinear problem all-at-once instead of solving a sequence of linear problems distinguishes our approach from the preconditioning strategy proposed in [18]. Moreover, although the model problem in [19] is very similar to ours, their ansatz for preconditioning is completely different from ours.
@@
In Sect. 2 we introduce the proposed method; first, we describe the underlying setting and the idea in general, and then we discuss this approach in more detail on behalf of our two concrete model problems in Sects. 2.3 and 2.4. Numerical examples illustrating the advantages of our method are given in Sect. 3, and a prototypical application to probabilistic inverse problems is presented in Sect. 4. We conclude the paper with a brief summary and outlook on future work in Sect. 5. Finally, A contains the proof of a fact that can be understood as a heuristic motivation for the chosen approach.
@@ 2.1 Problem statement
Let now $V_{h}\subset V$ be a finite dimensional subspace, e.g., resulting from a conforming finite element discretization of $V$; the index $h$ indicates the discretization parameter, e.g., the mesh size used for finite element discretization. The so-called Galerkin discretization of our problem (2) is given by:
@@ 2.2 Proposed Method
Since (4) is a possibly nonlinear system of equations, usually consisting of a high number of variables, solving (4) for several choices of $\lambda$ may become costly. To deal with this issue, we propose a method that makes use of the strong approximation capabilities of neural networks. More precisely, we want to approximate the map $\lambda\mapsto{\bf u}_{h}(\lambda)$ solving (4) by a neural network ${\bf u}_{h,\textnormal{NN}}^{\theta}\!\colon\mathbb{R}^{p}\rightarrow\mathbb{R}^{N_{h}}$; hereby, $\theta$ denotes the network parameters of the neural network. Training this neural network — or in other words: finding the right parameters $\theta$ that make ${\bf u}_{h,\textnormal{NN}}^{\theta}$ a hopefully good approximation of $\lambda\mapsto{\bf u}_{h}(\lambda)$ — requires to determine a suitable so-called loss function. Previous work [28] has suggested to use the $\ell^{2}$-norm of the discrete equation residual, i.e., to determine $\theta=\operatorname{argmin}_{\vartheta}\mathcal{L}$ of
@@
with ${\bf P}^{L}_{\lambda,h},{\bf P}^{R}_{\lambda,h}\in\mathbb{R}^{N_{h}\times N_{h}}$ being the left, and respectively right, preconditioners. The desired ${\bf u}_{h,\textnormal{NN}}^{\theta}(\lambda)$ can then be reobtained via ${\bf u}_{h,\textnormal{NN}}^{\theta}(\lambda)={\bf P}^{R}_{\lambda,h}{\bf\tilde{u}}_{h,\textnormal{NN}}^{\theta}(\lambda)$. Of course, such a preconditioner must be chosen carefully taking into account the problem characteristics. As already mentioned in the introduction, we directly apply preconditioning the the possibly nonlinear residual, which is in contrast to [18] where nonlinear problems are treated by solving linear problems in iterative manner. Moreover, unlike in [18] we precondition from both sides.
@@
To optimize the loss function w.r.t. the neural network parameters $\theta$, usually variants of the gradient descent method, such as the Adam optimizer [36], or quasi-Newton methods, such as the limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) algorithm [37] are employed. These methods require the gradient of the loss function with respect to the network parameters, which is calculated by using automatic differentiation (AD) [9]; thus all computations in the FEM Module, such as, e.g., setting up stiffness and/or mass matrices etc., must be implemented automatically differentiable. In this paper we will solely employ L-BFGS. Like most optimizers popular in deep learning, L-BFGS only requires gradients of the objective function, i.e., first-order information, and thus avoids the expensive computation of second-order derivatives by AD, which would be necessary for second-order methods as, e.g., Newton’s method. Nevertheless, as a quasi-Newton method, L-BFGS lies somewhere between first- and second-order methods as it is able to exploit curvature information by successively building an approximation to the inverse Hessian matrix during the iteration. The associated hopes for faster convergence have been confirmed, e.g., in the case of classic PINNs [8].
@@
Hence, we hope to improve the conditioning of the Hessian of $\mathcal{L}_{{\bf P}_{h}}$ at some minimizer $\theta$ by choosing ${\bf P}^{L}_{\lambda,h}$, ${\bf P}^{R}_{\lambda,h}$ in such a way that the condition number of $\mathbf{P}^{L}_{\lambda,h}\cdot[D_{\bf u}{\bf R}_{\lambda,h}](\mathbf{P}^{R}_{\lambda,h}{\bf\tilde{u}}_{h,\textnormal{NN}}^{\theta}(\lambda))\cdot\mathbf{P}^{R}_{\lambda,h}$ becomes relatively small. Even if these considerations by no means constitute a strict proof, we consider them to be a conclusive motivation for our approach, and our numerical observations in Sect. 3 seem to support this.
@@ 2.3 Example 1: Stokes equations in 2D
In the following, we show some more concrete details regarding the realization and implementation of the proposed FEM-based neural network for solving the Stokes flow around an airfoil with a parameterizable angle of attack. We consider the problem in a 2-dimensional Lipschitz-bounded domain $\Omega$, where the Stokes problem consists of finding two functions, the velocity $u(\lambda)\!:\Omega\rightarrow\mathbb{R}^{2}$ and the pressure $p(\lambda)\!:\Omega\rightarrow\mathbb{R}$ such that
@@
with functions $f_{\lambda}\!:\Omega\rightarrow\mathbb{R}^{2}$, $g_{\lambda}\!:\Omega\rightarrow\mathbb{R}$ and suitable boundary conditions, which will be introduced precisely below. Equation (7) represents the conservation of momentum, Equation (8) the conservation of mass. Of course, depending on the regularity of the formulation one seeks, these equations can be understood in a strong sense, almost everywhere, or, if $f,g$ become distributions, in a distributional sense.
For the sake of simplicity we will focus in the following on $f\in L^{2}(\Omega)^{2}$ and $g\in L^{2}(\Omega)$, and choose function spaces $X=\{x\in H^{1}(\Omega)^{2}\mathrel{}\mathclose{}|\mathopen{}\mathrel{}\left.x\right|_{\partial\Omega_{D}}=0\}$ for velocity and $M=L^{2}(\Omega)$ for pressure. Here, $\partial\Omega_{D}$ is supposed to signify the Dirichlet boundary, which shall not comprise the entire boundary $\partial\Omega$. For the time being, we impose zero as the Dirichlet boundary condition for the velocity, which will be alleviated later on.
A weak formulation of the system (7)–(8) with a conforming Galerkin ansatz space $V_{h}:=X_{h}\times M_{h}\subset X\times M$ then reads as follows:
@@
with the bilinear forms $a(u_{h},v_{h}):=\int_{\Omega}\nabla u_{h}:\nabla v_{h}$ and $b(v_{h},p_{h}):=-\int_{\Omega}p_{h}\nabla\cdot v_{h}$. The two functionals $f$ and $g$ are defined as $f(v_{h})\coloneqq\int_{\Omega}v_{h}f_{\lambda}$ and $g(q_{h})\coloneqq-\int_{\Omega}q_{h}g_{\lambda}$, respectively. $R_{\lambda}$ is accordingly defined as the residual of Equation (9). Note that the fully continuous variant, that is, setting $X_{h}=X$ and $M_{h}=M$, defines a well-posed problem [39, Prop. 4.7]. However, to obtain a well-posed problem in the discrete setting, the discrete problem must satisfy the Ladyzhenskaya–Babus̆ka–Brezzi (LBB) condition [40, 41, 42] and hence, the choice of $X_{h},M_{h}$ must be handled with care. We use Taylor–Hood mixed finite elements, i.e., continuous, piecewise linear ansatz functions for the pressure and continuous, piecewise quadratic ansatz functions for the velocity. With this choice of $V_{h}$, the LBB condition for (9) is satisfied and we obtain a well-posed discrete problem  [39, Lemma 4.23].
Figure 2: Coarse triangular mesh around the NACA 0012 airfoil.
@@
Lemma 2.1 in [44] shows that the conditioning of the plain saddle point problem (10) is bounded from below by the conditioning of the discretization of the Laplacian. As such, classical FEM theory [39, Theorem 9.14] shows that for a Lagrange finite element, as we have chosen for the velocity discretization, the conditioning scales with $h^{-2}$. Therefore, it is usually necessary to apply preconditioning to the simple saddle-point problem. In fact, the condition number of the square of the preconditioned stiffness matrix in (12) is bounded independently of the discretization, as shown in A. There, it is even shown that the number of distinct eigenvalues of this matrix is three. Considering again that BFGS with perfect line search is equivalent to CG for quadratic optimization problems [38], BFGS applied to such an quadratic optimization problem with the squared preconditioned matrix as its Hessian, would converge in at most three iterations [^2] The situation is analogous to the Schur-preconditioned stable FEM solutions of the Stokes problem, when solved with Krylov subspace methods, such as MINRES, cf. [45, p. 292].. Of course, these properties do not directly transfer to our setting, as the problem under consideration is highly non-linear and non-convex due to the presence of a neural network. Nevertheless, it does not seem unreasonable to assume that applying the preconditioner to the linear part of the problem will lead to at least a gradual improvement of the overall properties in this case as well. In fact, our numerical examples in Sect. 3 support this.
@@
Let us now turn to the explicit numerical example, where we want to solve the Stokes equations (7)-(8) with source terms $f_{\lambda}=0,g_{\lambda}=0$, and non-homogeneous Dirichlet boundary conditions. We define the domain $\Omega$ as $(0,5)^{2}$ with a NACA 0012 airfoil [46] of length one cut out in the center (see Figure 2). A Dirichlet boundary condition dependent on the angle of attack $\lambda$ is used at the inflow boundary $\partial\Omega^{+}:=\{(x,y)\;|\;x=0\;\text{or}\;y=0\}$, a Neumann boundary condition at the outflow boundary $\partial\Omega^{-}:=\{(x,y)\;|\;x=5\;\text{or}\;y=5\}$ and a no-slip boundary condition at the airfoil boundary $\partial\Omega^{A}$:
$u$ $=[\cos(\lambda\pi/180),\textnormal{ }\sin(\lambda\pi/180)]^{T}$ $\;\text{on}\;\partial\Omega^{+},$
$p-\partial_{\nu}u$ $=0$ $\;\text{on}\;\partial\Omega^{-},$
$u$ $=0$ $\;\text{on}\;\partial\Omega^{A}.$
Here $\partial_{\nu}$ signifies the partial derivative in normal direction on the boundary. Note that the homogeneous Neumann boundary condition on the outflow boundary effectively sets the mean outflow pressure to zero.
Homogeneous Neumann boundary conditions are known to be so-called natural conditions and thus included in (10) by construction, and do not need any special treatment. However, our aforementioned theory assumes homogeneous Dirichlet boundary conditions for the velocity as well. In order to enforce the inhomogeneous boundary conditions, we will implicitly construct a lifting function: Firstly, we assemble all the matrices and residuals oblivious to the Dirichlet boundary conditions. Secondly, we exploit that the chosen basis functions are Lagrangian, i.e., each basis function $v_{h}^{i}$ has an associated Lagrangian node $x_{i}\in\partial\Omega$ such that
$v_{h}^{i}(x_{j})=\delta_{ij}\quad\forall i,j.$ (13)
The Dirichlet boundary conditions are now enforced on all such $x_{j}\in\partial\Omega^{+}\cup\partial\Omega^{A}$ by adjusting the right-hand side vector $\mathbf{f}$ and the stiffness matrix in (10) accordingly: First, we multiply the columns of the stiffness matrix corresponding to the boundary nodes by the specified boundary values and then subtract the result from $\mathbf{f}$. Next, we set the corresponding values of $\mathbf{f}$ to the boundary values. In the stiffness matrix we modify the rows and columns corresponding to the boundary nodes so that the diagonal value is unity and the off-diagonal entries are set to zero. Note that this is done before preconditioning, which is possible because the modifications do not affect the symmetry properties of the system (10).
It is also noteworthy that, by the aforementioned approach of incorporating the Dirichlet boundary conditions, we obtain in principle a single objective, avoiding the issues of weakly imposing the Dirichlet conditions via an additional objective term. However, we believe that avoiding the multi-objective scenario is not per se an issue, as one could also rewrite our objective with multiple terms. In our view, the more important part is that these potentially additional terms that impose the Dirichlet conditions must be mutually non-conflicting. We explicitly ensure this by erasing parts of the assembled stiffness matrix and adequately modifying the source terms, as mentioned above.
2.4 Example 2: Stationary Navier–Stokes equations in 2D